OpenStudy (anonymous):

The average height of an apple tree is 4.00 meters. How long would it take an apple falling from that height to reach the ground? Given: g = -9.8 meters/second2 A.) 0.90 sec. B.) 1.0 sec. C.) 1.53 sec. D.) 2.3 sec.

OpenStudy (anonymous):

well the dissplacement which is 4m can be given by s=1/2gt^2 now lets insert the known values into this ignoreing the (-) 4m=1/2(9.8xt^2) all you need to do now is change the subject of the formulea to slove for time.\[\sqrt{\frac{ 2s }{ g }}=t\] and you con calculate the rest. answer should be (A)

OpenStudy (anonymous):

thank you! @zakisista

OpenStudy (anonymous):

your welcome, if you do not fully understand this then please ask me, i would love to help.